OpenStudy (anonymous):

could someone help me find antiderivaative of 5/x^4

OpenStudy (agreene):

\[\int \frac{5}{x^4}\space dx=-\frac{5}{3x^3}+C\]

OpenStudy (anonymous):

know that 5/x^4=5x^(-4)

OpenStudy (anonymous):

could you explain the steps to find that

OpenStudy (agreene):

\[\int\frac{5}{x^4}dx=5\int\frac{1}{x^4}dx=5\int x^{-4}dx=5(3x^{-3})+C=\frac{5}{3x^3}+C\]

OpenStudy (anonymous):

hmm...okay i think i kinda see it

OpenStudy (agreene):

I left off a negative sign when I took the integral of x^-4...

OpenStudy (agreene):

should be 5 \integral x^-4 dx = 5(-3x^-3) = - 5/(3x^3)+C I switched to the negative exponents because it makes it a bit easier, like doing a regular exponent... I don't generally do that so I tend to mess it up.